Suppose a country has total GDP(Y)=\$12 trillionconsumption (C)=\$8 trillion government spending (G)=\$2 trillion Investment $3 trillionand taxes . a. What is the level of net exports or balance of tra $trillion bWhat is the level of public savings? What is the level of private savings? $ trillion What is the level of net capital outflow ?


1
Expert's answer
2021-12-12T17:53:20-0500

a) Net exports:


Net exports can be calculated as:


GDP= consumption + Investment + Government expenditure + NEt export


12 = 8+3+2+NEt export


Net export = -1


Hence, the value of net export is -$1 trillion




(b) Public savings


Public savings can be calculated as follows:


Public savings = Tax - Government expenditure


=3 - 2


=1


Hence, the value of public saving is $1 trillion




(c) Private savings:


Private savings can be calculated as follows:


Private saving= (GDP - Tax) - Consumption


=(12 - 3) - 8


= 1


Hence, the value of public savings is $1 trillion




(d) Net capital flow:


Net capital flow can be calculated as follows:


National capital flow= Saving - Investment


=(GPD - Consumption - government expenditure) - Investment


= (12- 8 - 2) - 3


= -1


Since the capital flow is negative, there is net capital outflow and is $1 trillion